OpenStudy (anonymous):

Solve: 1=(3)/(x+3)+(18)/(x-3)^(2) Thx ;)

OpenStudy (anonymous):

\[\frac{3}{x+3}+\frac{18}{(x-3)^2}=1\] \[\frac{3(x-3)^2+18(x+3)}{(x+3)(x-3)^2}=1\] \[3(x^2-6x+9)+18x+54=(x+3)(x-3)^2\] \[3x^2-18x+27+18x+54=(x+3)(x^2-6x+9)\] \[3x^2+81=x^3-6x^2+9x+3x^2-18x+27\] \[x^3-6x^2-9x-54=0\]

OpenStudy (anonymous):

sure it wasn't (3)/(x-3)

OpenStudy (anonymous):

Ooh waoo, Thx.. :)

OpenStudy (anonymous):

its not solved yet...

OpenStudy (anonymous):

you sure you typed your question right?

OpenStudy (anonymous):

it is not solved...somenthing is wrong ://

OpenStudy (anonymous):

Ohh, I am so sorry... its x+3 on both sides..

OpenStudy (anonymous):

that makes it much easier...you should make sure you type your question right if you want the right answer

OpenStudy (anonymous):

\[\frac{3}{x+3}+\frac{18}{(x+3)^2}=1\] This is your problem?

OpenStudy (anonymous):

Yes, Thx again.

OpenStudy (anonymous):

(x+3)^2 is the LCD \[\frac{3(x+3)+18}{(x+3)^2}=1\] \[3x+9+18=(x+3)^2\] \[3x+27=(x+3)^2\] \[3x+27=x^2+6x+9\] \[x^2+3x-18=0\] \[(x+6)(x-3)=0\] x=-6, 3

OpenStudy (anonymous):

Thumb up!! Thx.

OpenStudy (anonymous):

3 ( x+3) + 18 = ( x+3) ^2 3x + 27 = x^2 + 6x + 9 -> x^2 + 3x - 18 = 0 =>( x + 6) ( x -3) = 0 Thus x = -6, x = 3